maureen paid $62.06 including 7% tax, for a sweater. How much did the sweater itself cost?
X=62.06-7/100
x= 62.06-.07
x=-61.99
this is not right could you help me solve?

Answer by gonzo(654)

x = cost of the sweater
x + .07*x = 62.06
1.07*x = 62.06
x = 62.06/1.07
x = 58
-----
58*1.07 = 62.06
-----